# MedUnifier: Unifying Vision-and-Language Pre-training on Medical Data with Vision Generation Task using Discrete Visual Representations

### v2

@@ 4.4 Ablation study
To demonstrate the efficacy of our proposed method, we perform an ablation study (Table 6) across various learning objectives. The results indicate that using only the ITC loss (ID 1) yields the lowest performance. As ITM and ITG objectives are incrementally incorporated, performance gradually improves (IDs 2 and 3), highlighting how refined cross-modal alignment and multi-modal language modelling enhance the model’s overall capabilities. Interestingly, the model with TIG (ID 4) surpasses the one with ITG (ID 3). We attribute this phenomenon to the relative difficulty of generating pixel-level image representations guided by text, as compared to generating word-level representations guided by visual input, which leads the model to learn more abstract, well-aligned representations. Ultimately, the integration of all objective types (ID 5) enables the model to achieve optimal performance, underscoring the viability of incorporating vision generation into existing frameworks. More elaborate ablation study over more downstream tasks can be found in appendix 5.
@@ 5 Conclusion
In this paper, we introduce a novel and unified Med-VLP model, MedUnifier, which optimizes four distinct learning objectives simultaneously. By leveraging learnable embeddings and encoding raw images through a pre-trained Vision Transformer (ViT), MedUnifier circumvents the need to learn visual embeddings from scratch. Additionally, a VQ-VAE-based text-grounded image generation task is further incorporated into the Med-VLP framework to enhance its representation learning capacity. It reconstructs pixel-level visual details from both image and report, facilitating fine-grained visual understanding commonly available in medical data (subtle visual details e.g. small nodules, slight opacities, etc.) and efficient use of multi-modal representations through hierarchical latent adapters of dynamically adjusting the abstraction for each mode. The adapters help align vision and texts, ensuring that the each data mode contributes its most relevant features. Our proposed method effectively complements existing Med-VLP frameworks and achieves state-of-the-art performance on various downstream tasks, as validated by experimental results. Our work also has significant implications for enhancing the VLP development for radiological applications.
@@ 6 Acknowledgement
This project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RG25/24). This project is supported by the Lee Kong Chian School of Medicine - Ministry of Education Start-Up Grant. This project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RS16/23). This research was supported in part through the computational resources and staff contributions provided for the Quest high performance computing facility at Northwestern University which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology.
Appendix
@@ 2 Implementation details
For the image-text encoder, we adopt the pre-trained ViT-g from Fang et al. (2023) as preliminary visual embeddings extractor ($E_{I}$) and its hidden dimension ($d_{v}$) is set to 1408. We adopt a base-size Transformer encoder with 12 layers initialized from Boecking et al. (2022) and its hidden dimension ($d_{q},d_{t}$) is set to 768. The length of learnable embeddings was set to 32 ($L_{q}$). The top and bottom codebook size were both set to 512 with feature dimension of 768. The top image decoder consisted of a deconvolutional block (2x upsample a 2D feature map) where there are a transposed convolution, a relu activation and a batch normalization, upsampling top latent encodings into bottom latent encodings. The bottom image decoder consisted of four deconvolutional blocks, decoding bottom latent encodings into original pixels. For tokenizing medical reports specialized medicine vocabulary dictionary from Boecking et al. (2022) was employed with vocabulary size of 30522. All text inputs were confined to a maximum length of 95 ($L_{t}$). All images size was resized into $3\times 224\times 224$ ($C\times H\times W$) and normalized into range $(0,1)$. Raw images were split into non-overlap patches by stride of $14\times 14$ ($h\times w$) so that patch embeddings length was $16*16=256$ ($L_{v}$). The feature dimension of ITC linear heads was set to 256. For the text generator, we set the dimension of the prediction head to 768. For the image encoder, the size of latent feature maps were set to be $8\times 8,16\times 16$ for top level and bottom level, respectively. For optimization, we adopt the AdamW optimizer Loshchilov (2017) with $\beta_{1}=0.9,\beta_{2}=0.95$, and a weight decay of 0.05. We used a cosine learning rate decay with a peak learning rate of 1e-4. We used the warm-up strategy during the first 5% of the total number of steps and an initial learning rate of 1e-5. The pre-training process was running on four 80G NVIDIA A100 GPUs. We used the mix precision wuth Accelerate open-source library Gugger et al. (2022) to speed up training and save computational costs.
@@
For uni-modal downstream tasks, we used the AdamW optimizer Loshchilov (2017) with the learning rate set to 3e-6 and 3e-4 for the pre-trained model and task-specific layers, respectively. We conducted binary classification and multi-class classification for different datasets. For image-text retrieval tasks, we computed pairwise similarity to rank paired data by relevance. For zero-shot categorization we used images as queries and generated expert text prompts as in Huang et al. (2021). The text prompt with the highest score would be considered a predictive positive sample. We used images as the prompt for the medical report generation task to guide text generation. We trained two auto-regressive models in image generation, e.g. Pixelsnail Chen et al. (2018), to model multi-modal priors. Then we sampled latent encodings and fed them into a hierarchical decoder to generate new images.
@@ 3 Datasets
MIMIC-CXRJohnson et al. (2019)
This is the largest radiology dataset currently available, comprising chest X-ray images and corresponding reports from Beth Israel Deaconess Medical Center. It contains over 370,000 images from more than 65,000 patients, making it one of the most extensive collections of de-identified chest X-rays available for research. Each image is accompanied by detailed textual reports, which provide diagnostic information and contextual clinical notes. For the use of this dataset, We exclude samples that lack a "findings" or "impression" section within their clinical reports and retain only images in the frontal view. For dataset splitting, we utilize the official train-test splits provided by MIMIC-CXR. For downstream tasks, following the approach in Cheng et al. (2023), we sample the MIMIC 5x200 subset and remove it from the training set to ensure robust evaluation.
@@
The original CheXpert dataset’s chest radiographs Irvin et al. (2019) are multi-labeled to accommodate numerous medical observations occurring at the same time. Because our zero-shot classification relies on identifying the most comparable target, having numerous alternative labels for a target can lead to results that are inconsistent across categories. As a result, following setting in Zhang et al. (2022); Huang et al. (2021), we employ CheXpert’s partial data to construct the CheXpert 5x200 dataset, which has 200 solely positive images for each of the CheXpert competition tasks: atelectasis, cardiomegaly, pneumonia, edema, and pleural effusion. In this dataset, each image has positive labels for only one condition.
RSNA PneumoniaShih et al. (2019)
The RSNA Pneumonia Detection Challenge dataset, developed by the Radiological Society of North America (RSNA), is a large, annotated collection of chest X-ray images specifically labelled for pneumonia detection. We use the stage 2 version. This dataset contains 30k frontal view chest radiographs labeled either as ”normal” or ”Pneumonia”. We sample raw 500 positives and 500 negatives for zero-shot classification. For fine-tuning, the train/valid/test split each constitutes 70%/15%/15% of the dataset, following Cheng et al. (2023).
COVIDx Wang et al. (2020)
It includes more than 30,000 CXR pictures from a global group of more than 16,600 patients. 16, 490 positive COVID-19 pictures from more than 2,800 patients are included in this collection. We make use of version 6 of this dataset. The task is to categorize each radiograph into three groups: normal, non-COVID pneumonia, and COVID-19. The data split follows Wang et al. (2022a).
SIIM-ACR PneumothoraxZawacki et al. (2019)
The SIIM-ACR Pneumothorax Segmentation Challenge is a collaborative machine-learning competition organized by the Society for Imaging Informatics in Medicine (SIIM) and the American College of Radiology (ACR). SIIM-ACR Pneumothorax contains 12954 X-ray chest images, together with image-level pneumothorax annotation and pixel-level segmentation mask if pneumothorax exists. We use them for downstream supervised classification as in Cheng et al. (2023).

### v3

@@ 4.4 Ablation study
To demonstrate the efficacy of our proposed method, we perform an ablation study (Table 6) across various learning objectives. The results indicate that using only the ITC loss (ID 1) yields the lowest performance. As ITM and ITG objectives are incrementally incorporated, performance gradually improves (IDs 2 and 3), highlighting how refined cross-modal alignment and multi-modal language modelling enhance the model’s overall capabilities. Interestingly, the model with TIG (ID 4) surpasses the one with ITG (ID 3). We attribute this phenomenon to the relative difficulty of generating pixel-level image representations guided by text, as compared to generating word-level representations guided by visual input, which leads the model to learn more abstract, well-aligned representations. Ultimately, the integration of all objective types (ID 5) enables the model to achieve optimal performance, underscoring the viability of incorporating vision generation into existing frameworks.
@@ 5 Conclusion
In this paper, we introduce a novel and unified Med-VLP model, MedUnifier, which optimizes four distinct learning objectives simultaneously. By leveraging learnable embeddings and encoding raw images through a pre-trained Vision Transformer (ViT), MedUnifier circumvents the need to learn visual embeddings from scratch. Additionally, a VQ-VAE-based text-grounded image generation task is further incorporated into the Med-VLP framework to enhance its representation learning capacity. It reconstructs pixel-level visual details from both image and report, facilitating fine-grained visual understanding commonly available in medical data (subtle visual details e.g. small nodules, slight opacities, etc.) and efficient use of multi-modal representations through hierarchical latent adapters of dynamically adjusting the abstraction for each mode. Our proposed method effectively complements existing Med-VLP frameworks and achieves state-of-the-art performance. Our work also has significant implications for enhancing the VLP development for radiological applications.
@@ 6 Acknowledgement
This project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RG25/24 and RS16/23), as well as the Lee Kong Chian School of Medicine - Ministry of Education Start-Up Grant. This research is also partially supported by the computational resources and staff contributions of the Quest high-performance computing facility at Northwestern University, jointly funded by the Office of the Provost, the Office for Research, and Northwestern University Information Technology.
Supplementary Material
@@ 8 Implementation details
For the image-text encoder, we adopt the pre-trained ViT-g from [18] as preliminary visual embeddings extractor ($E_{I}$) and its hidden dimension ($d_{v}$) is set to 1408. We adopt a base-size Transformer encoder with 12 layers initialized from [4] and its hidden dimension ($d_{q},d_{t}$) is set to 768. The length of learnable embeddings was set to 32 ($L_{q}$). For tokenizing medical reports specialized medicine vocabulary dictionary from [4] was employed with vocabulary size of 30522. All text inputs were confined to a maximum length of 95 ($L_{t}$). All images size was resized into $3\times 224\times 224$ ($C\times H\times W$) and normalized into range $(0,1)$. Raw images were split into non-overlap patches by stride of $14\times 14$ ($h\times w$) so that patch embeddings length was $16*16=256$ ($L_{v}$). The feature dimension of ITC linear heads was set to 256. For the text generator, we set the dimension of the prediction head to 768. For the image encoder, the size of latent feature maps were set to be $8\times 8,16\times 16$ for top level and bottom level, respectively. For optimization, we adopt the AdamW optimizer [43] with $\beta_{1}=0.9,\beta_{2}=0.95$, and a weight decay of 0.05. We used a cosine learning rate decay with a peak learning rate of 1e-4. We used the warm-up strategy during the first 5% of the total number of steps and an initial learning rate of 1e-5. The pre-training process was running on four 80G NVIDIA A100 GPUs. We used the mix precision wuth Accelerate open-source library [21] to speed up training and save computational costs.
@@
For uni-modal downstream tasks, we used the AdamW optimizer [43] with the learning rate set to 3e-6 and 3e-4 for the pre-trained model and task-specific layers, respectively. We conducted binary classification and multi-class classification for different datasets. For image-text retrieval tasks, we computed pairwise similarity to rank paired data by relevance. For zero-shot categorization we used images as queries and generated expert text prompts as in [26]. The text prompt with the highest score would be considered a predictive positive sample. We used images as the prompt for the medical report generation task to guide text generation. We trained two auto-regressive models in image generation, e.g. Pixelsnail [8], to model multi-modal priors. Then we sampled latent encodings and fed them into a hierarchical decoder to generate new images.
@@ 9 Datasets
MIMIC-CXR[30]
This is the largest radiology dataset currently available, comprising chest X-ray images and corresponding reports from Beth Israel Deaconess Medical Center. It contains over 370,000 images from more than 65,000 patients, making it one of the most extensive collections of de-identified chest X-rays available for research. Each image is accompanied by detailed textual reports, which provide diagnostic information and contextual clinical notes. For the use of this dataset, We exclude samples that lack a ”findings” or ”impression” section within their clinical reports and retain only images in the frontal view. For dataset splitting, we utilize the official train-test splits provided by MIMIC-CXR. For downstream tasks, following the approach in [11], we sample the MIMIC 5x200 subset and remove it from the training set to ensure robust evaluation.
@@
The original CheXpert dataset’s chest radiographs [28] are multi-labeled to accommodate numerous medical observations occurring at the same time. Because our zero-shot classification relies on identifying the most comparable target, having numerous alternative labels for a target can lead to results that are inconsistent across categories. As a result, following setting in [69, 26], we employ CheXpert’s partial data to construct the CheXpert 5x200 dataset, which has 200 solely positive images for each of the CheXpert competition tasks: atelectasis, cardiomegaly, pneumonia, edema, and pleural effusion. In this dataset, each image has positive labels for only one condition.
RSNA Pneumonia[53]
The RSNA Pneumonia Detection Challenge dataset, developed by the Radiological Society of North America (RSNA), is a large, annotated collection of chest X-ray images specifically labelled for pneumonia detection. We use the stage 2 version. This dataset contains 30k frontal view chest radiographs labeled either as ”normal” or ”Pneumonia”. We sample raw 500 positives and 500 negatives for zero-shot classification. For fine-tuning, the train/valid/test split each constitutes 70%/15%/15% of the dataset, following [11].
COVIDx [59]
It includes more than 30,000 CXR pictures from a global group of more than 16,600 patients. 16, 490 positive COVID-19 pictures from more than 2,800 patients are included in this collection. We make use of version 6 of this dataset. The task is to categorize each radiograph into three groups: normal, non-COVID pneumonia, and COVID-19. The data split follows [58].
SIIM-ACR Pneumothorax[66]
The SIIM-ACR Pneumothorax Segmentation Challenge is a collaborative machine-learning competition organized by the Society for Imaging Informatics in Medicine (SIIM) and the American College of Radiology (ACR). SIIM-ACR Pneumothorax contains 12954 X-ray chest images, together with image-level pneumothorax annotation and pixel-level segmentation mask if pneumothorax exists. We use them for downstream supervised classification as in [11].